a crate is pushed along a floor with a force of 450 N. if frictional force is 120 N and gravitational force is 4000 N what is the net force of the crate?
1
Expert's answer
2020-12-15T11:33:41-0500
"F_{net}=F-F_f\\F_{net}=450-120=330\ N"